29 September 2011: MODS checkout/OSURC science

afternoon Loaded new MOS masks for AZ run and reshuffled permanent masks. See the MODS Mask Exchange Procedure for details of the day's activities.

01:29 cd ~MODSeng/modsScripts/Support/
execMODS modsWake.pro
Observing mode
Geno restarts GCS

01:30 need to reload slit table (Rick did this for the permanent masks, not yet for the User masks).

02:08 Collimated on Persson star. Must do initial (or after reopening) collimation on-axis because it takes so long. Once RMS WFE ~400nm, Slewed to 1717. Cloud just came over the field... Aborted failed preset (failed because no star found).

02:17 Aborted (with x the script --- so I did not take the confirmatory image). The guide star was lost due to clouds. 02:40 closed for thick clouds.

04:10 Re-opened. Re-checked pointing and collimation.

04:30 acqMODS mkn957.acq 3 was slit and 4 was field image, 5 was confirmatory image.

04:40 approx execMODS mkn957.obs dual 0.6" slit 2 x 600 sec

05:05 acqMODS feige110.acq (NB: 20 sec is too long for acq on this star).

05:12 execMODS feige110.obs

05:19 acqMODS phl1092.acq (clouds came in during acquisition and guide star was lost).

05:36 resent preset and acquired

05:41 execMODS phl1092.obs dual 0.6" slit 2 x 1350 sec.
  • Stopped 2nd exposure about 1/2 way through (and read it out) because of clouds at 06:20

06:23 acqMODS UM_RHMstarsM33C-4119E_ls.acq (E of meridian. PI has set up scripts to use PA ~ PARANGLE)

06:32 execMODS UM_RMHstarsM33C-4119_ls.obs 3 x 240 sec. dual 0.6" slit
  • 0.6" fwhm measured from acq image. Guide star fwhm ~ 0.9".

06:50 execMODS OSU_LSNeIm_SN1995av.img imaging script using g and z filters 1K x 1K
  • 06:57 clouds come in ... guide star fading in and out.

07:12 execMODS OSU_LSNeIM_SN2005gj25as.img --- canceled because guide star was not found ---

07:40 back to SN1995av. 11th magnitude guide star comes and goes... ugh!

08:43 acqMODS mkn1044.acq red 40 is slit, 41 is field

08:50 execMODS mkn1044.obs 0.6" dual grating

09:06 acqMODS iras03450.acq (bypassed confirmatory image with 'x' from acq script)

09:14 execMODS iras03450.obs 0.6" dual grating
  • there are clouds. This script does 2 x 900 sec integrations.
  • blue image 26 shows some artifacts of bright V~8 star from previous sn1995av imaging. The horizontal register was blasted by this star... it takes a few erase cycles to get rid of charge.

09:49 acqMODS iras04416.acq ... guide star signal too faint with clouds

09:52 acqMODS OSU_LSNeSpec_2000ei.acq
  • difficult acquisition. Manually took exposure 2x longer, but barely see target in r image.

10:04 execMODS OSU_LSNeSpec_2000ei.acq
  • seeing is blowing up to 2.00" now.
  • obs script set up for 12 x 1200 seconds, but we'll see what we get after a single 1200 seconds exposure.
  • stopped during 2nd exposure, with 900 seconds remaining b 29 and r 53

10:34 acqMODS iras04416.acq (again, now that transparency is better).
  • Guide star selected was fainter than 17th mag! Rebecca is editing script to change guide star. This requires change of PA from 0 to 180 deg. New guide star is ~ 14th mag. It is double (not seen on dss image) but with seeing 1.2" they seem to be well separated and not affecting guiding.
  • 56 and 57 are slit and field images, 58 is thru-slit confirmatory image.

10:48 execMODS iras04416.obs 0.6" dual grating 2 x 600 sec exposures.

11:13 acqMODS g191b2b.acq (Rebecca edited script, in OSU_FeIIAGN directory, to take 20 sec slit image, as originally written, but only 0.5 sec field image, because even 2 seconds saturated on feige110 (V=11.5), only 0.3 mag brighter than G191B2B and the same B-V color.) On field image (red 62), through r-sdss, star has about 7000 counts in peak.

11:21 execMODS g191b2b.obs. 0.6" dual 4 x 30 seconds. Seeing is again back to about 1" on guider.
  • 30 sec spectrum has about 2000 counts.

11:30 acqMODS iras07598.acq At 45deg elevation and 65 deg N, collimation was initially pretty far out... There are 1-2 rows of lenslets on the left that are not illuminated. This is after about 6-8 cycles. Taking a 2nd field image, the first one showed stars with tails, about 19 pixels long. 2nd acquisition.
  • 1st acq: 69 tails extending to the upper right about 19 pixels
  • 2nd acq: 70 (WFE about 600) mods1r.20110929.0070.fits
  • 3rd acq 71 (done when WFE was 470 or lower). This one looks better.
  • 72 was thru-slit confirmatory image

11:45 execMODS iras07598.obs

12:20 script ended and it was getting light.
  • put MODS in calibration mode (oops! forgot to stop guiding... guider complained.)
  • Geno closing dome and parking rotator at 210 degrees.
  • Back to zenith. Dome lights off
  • Sept29.multcal is started at 05:31
    • 05:39 blue-only ended, red-only ended 05:44, dual ended
    • imaging flats were aborted (ERROR: No operating mode has been set in this script. (must have OBSMODE, CALMODE or SETUPMODE).
  • Sept29.multcal (without the imaging flats) finished at 06:21 am.
  • ran modsWake
  • ran Sept29.imf.multcal to do the imaging flats and put the instrument to sleep. (error - forgot slitmask imaging...add it and run).

#execMODS -u bgrflats_0.8.cal
#execMODS -u grslitflat.cal
#execMODS -u grflats.cal
#execMODS -u bgrflats.cal
#execMODS -u rgrflats.cal
execMODS -u bgrlamps.cal
execMODS -u rgrlamps.cal
execMODS -u grlamps.cal
execMODS -u blueimflat.cal
execMODS -u redimflat.cal
execMODS -u /home/MODSeng/modsScripts/Support/modsSleep.pro
execMODS -u bias1K.cal
execMODS -u bias3K.cal
execMODS -u bias8K.cal

execMODS -u blueimflat.cal
execMODS -u redimflat.cal
execMODS -u /home/MODSeng/modsScripts/Support/modsSleep.pro

Blue g_sdss flats were OK, 40K, Vflat 4.0 and 20 sec. Blue u_sdss flats were saturated in bottom right quadrant QTH2 and 1 sec. Red r_sdss flats Vflat 2.0 and 25 sec OK, 20K Red i_sdss flats Vflat 1.5 and 15 sec OK, 40K Red z_sdss flats Vflat 1.5 and 8 sec a bit high, 55-57K

-- OlgaKuhn - 29 Sep 2011
Topic revision: r10 - 30 Sep 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback